OpenStudy (anonymous):

To go on the class trip, Chang needs at least $125. He has saved $45, and he makes $7 an hour at his job. What is the minimum number of complete hours he must work to have enough money to go on the trip? (Assume there are no withholdings from his check.) 10 hours 11 hours 12 hours

OpenStudy (callisto):

Let x be the hours he needs to work. \[45+7x\ge 125\] Can you solve it

OpenStudy (anonymous):

ill try

OpenStudy (callisto):

Note: 7x is the amount he earns through working 45 is the money he saved 125 is the sum of money he needed

OpenStudy (anonymous):

12 because its more

OpenStudy (callisto):

Can you show all your workings?

OpenStudy (anonymous):

45+7(12)=129

OpenStudy (callisto):

Hmm... I don't think should do in this way :| Actually, it is like this: \[45 +7x \ge 125\]\[7x \ge 125 -45\]\[7x \ge 85\]\[7x \ge 80\]\[x \ge 11.429\] Therefore, the minimum hours required is 12